Consider a recurrence relation an = an-1 - 3an-2 for n = 1,2,3,4,… with initial conditions a1 = 3 and a2 = 5. Calculate a5.


1
Expert's answer
2021-11-26T05:14:47-0500

Let us consider a recurrence relation "a_n = a_{n-1} - 3a_{n-2}" with initial conditions "a_1 = 3" and "a_2 = 5." Let us calculate "a_5."

It follows that "a_3 = a_{2} - 3a_{1}=5-3\cdot 3=-4." Then "a_4 = a_{3} - 3a_{2}=-4-3\cdot5=-19."

We conclude that "a_5 = a_{4} - 3a_{3}=-19-3(-4)=-7."